Anonymous1babes:

What is the solution for the equation 6x − 8 = 4x? x = . (Input whole number only). (5 points) Blank 1:

AZ:

6x - 8 = 4x To solve for x, we have to get it all by itself on one side. What do you get if you subtract 6x on both sides?

Anonymous1babes:

0 for the left and -1 for the other

AZ:

No

AZ:

What is 6x - 6x = ? What is 4x - 6x = ?

Anonymous1babes:

x

AZ:

No

Anonymous1babes:

or is it 12x

Anonymous1babes:

idk how to to do this

AZ:

Imagine the 'x' as an apple You have 6 apples. You take away 6 apples. How many apples do you have left?

Anonymous1babes:

none

AZ:

Exactly, so 6x - 6x = 0 0x is the same thing as 0 zero is 0 is zero

AZ:

Now if you have 4 apples and you give away 6, how many are you left with? Now you're dealing with negative numbers here

Anonymous1babes:

thats what i said in the beginning i said 0 u said "no"

AZ:

No... it's not the same When you initially said 0, it makes no sense because there's still a -8 on the left hand side of the equal sign

Anonymous1babes:

ohhh

AZ:

Can you tell me what 4x - 6x equals?

Anonymous1babes:

-24

AZ:

We are not multiplying the numbers... Also don't forget that there should be an 'x' somewhere

AZ:

You need to subtract the numbers

Anonymous1babes:

ok

AZ:

and?

AZ:

if you don't know how to subtract 6 from 4... use a calculator

Anonymous1babes:

-2x

AZ:

Yes

AZ:

6x - 8 = 4x We subtracted 6x from both sides 6x - 8 - 6 = 4x - 6x Simplifying it now -8 = -2x We need to do one last thing to get 'x' all by itelf. Since we're multiplying the x by -2, to isolate the 'x' we need to do the opposite of multiplication. We need to divide it by -2. What happens when you divide both sides by -2? -8 = -2x

AZ:

What is -8 / -2 = ?

Anonymous1babes:

= 4?

AZ:

The negatives cancel out so essentially you're left with 8/2 = ?

AZ:

That is your answer!

AZ:

x = 4

Anonymous1babes:

thank u so much im not that good in math but at least u helped me through it :)

AZ:

It was my pleasure :)